OpenStudy (anonymous):

simplify : 1-(cos^2x/cot^2x)

OpenStudy (danjs):

Do you remember, the cotangent is the cosine over the sine

OpenStudy (danjs):

\[1 - \frac{ \cos^2(x) }{ \cos^2(x)/\sin^2(x) }\]

OpenStudy (danjs):

same as... \[1 - \frac{ \cos^2(x) }{ 1 }*\frac{ \sin^2(x) }{ \cos^2(x) }\]

OpenStudy (anonymous):

1-sin^2=cos^2x

OpenStudy (anonymous):

Is it right?

OpenStudy (danjs):

yeah , that is an identity

OpenStudy (anonymous):

so answer is cos^2x

OpenStudy (danjs):

yep, i think so

OpenStudy (danjs):

you could always try an easy angle in both of them and see if they come out the same

OpenStudy (anonymous):

Thank you so much for the help

OpenStudy (danjs):

welcome